A bag of colored blocks contains the following assortment of colors:
red (11), blue (22), orange (18), purple (13), green (10), and yellow (6).

Construct the probability distribution for x.

Answer by Positive_EV(69)

The probability distribution is a list of each possible outcome and its respective probability. There are six outcomes possible: red, blue, orange, purple, green, and yellow. The probability of drawing a block of a given color is the number of blocks of that color divided by the number of blocks total, assuming a block is drawn at random and all blocks have an equal chance of being picked. The total number of blocks is:

11 + 22 + 18 + 13 + 10 + 6 = 80

So, the probability of drawing a red block is 11/80, blue is 22/80 = 11/40, orange = 18/80 = 9/40, purple = 13/80, green = 10/80 = 1/8, and yellow = 6/80 = 3/40. This is the distribution of the probabilities of each outcome:

Red: 11/80
Blue: 11/40
Orange: 9/40
Purple: 13/80
Green: 1/8
Yellow: 3/40